Mr Rashid is 4 time as old his son.Four years ago,the sum of their ages was 37 years. Find the son's present age

Answer by FrankM(1040)

D = Dad's age now
S= Son's age now
D = 4 S
Four years ago, Dad was D-4 years old and the son was S-4
(D-4)+(S-4) = 37
D+S-8 = 37
D+S=45
But we know D=4S so substitute
4S+S = 45
S= 9 (and D= 36 in case you wonder)